4.A certain barangay is planning to implement a "quit smoking" program. In preparation, a survey was conducted to a sample of 200 smoking individuals asking who would like to join the program. One hundred and eighteen (118) expressed willingness to join the program. What are the values of p and q?

5.A total of 200 Grade 8 students who have access to Internet services were asked if they play games online before they attend classes. Ninety-five (95) students responded yes. What is the population proportion of students who play video games before attending classes? What percent do not play video games before attending classes?


1
Expert's answer
2021-03-30T09:38:49-0400

4. If 118 of 200 smoking individuals expressed willingness to join the program, then the values of p and q are:

p = 118/200 = 0.59 or 59%,

q = 1 - p = 1 - 0.59 = 0.41 or 41%.

The population proportion of students who play video games before attending classes is: 95/200 = 0.475 or 47.5%.

The percent of students who do not play video games before attending classes is 1 - 0.475 = 0.525 or 52.5%.